OpenStudy (anonymous):

what is equivalent to (11) / ((radical 3)-5)? can you explain how you did it?

OpenStudy (anonymous):

\[\frac{ 11 }{ \sqrt{3}-5 }\]

OpenStudy (anonymous):

thats what i mean ^

OpenStudy (anonymous):

To simplify (write what's equivalent), you have mutliply both the top and bottom by the conjugate For example if you had \[\Large \sqrt{a}+x\] then the conjugate would be \[\Large \sqrt{a}-x\]

OpenStudy (anonymous):

Therefore, we now have \[\Large \frac{11}{\sqrt{3}-5} \times \frac{\sqrt{3}+5}{\sqrt{3}+5}\]

OpenStudy (anonymous):

ok i get that please continue im still a little confused

OpenStudy (anonymous):

\[\Large \frac{11}{\sqrt{3}-5} \times \frac{\sqrt{3}+5}{\sqrt{3}+5}=\frac{11 \times (\sqrt{3}+5)}{\sqrt{3}^{2}-5^{2}}\]

OpenStudy (anonymous):

\[\Large \frac{11\sqrt{3}+55}{3-25}\] \[\Large \frac{11\sqrt{3}+55}{-22}\]

OpenStudy (anonymous):

When I did the denominator: \[\Large \sqrt{3}-5 \times \sqrt{3}+5\] I just remembered a rule that \[\Huge (a-b)^2=(a+b)(a-b)\] So it follows that

OpenStudy (anonymous):

\[\Large \sqrt{3}-5 \times \sqrt{3}+5=\sqrt{3}^2-5^2\] where a=sqrt(3) b=5

OpenStudy (anonymous):

ok i understand that but is there a way the simplify \[\frac{ 11\sqrt{3}+55 }{ -22 }\] any more?

OpenStudy (anonymous):

Sorry the rule is \[\Huge a^2-b^2=(a+b)(a-b)\]

OpenStudy (anonymous):

Yes, there is

OpenStudy (anonymous):

Factor out 11 in the top and the bottom.

OpenStudy (anonymous):

\[\frac{ 11\sqrt{3}+55 }{ -22 }=\frac{ 11*(\sqrt{3}+5 )}{11* (-2)}=-\frac{ \sqrt{3}+55 }{ 2 }\]

OpenStudy (anonymous):

oh ok thank you so much